UT 20130202 Technical Observing

Observer: JHill, MDeLaPena, CBiddick, DSummers, DMiller (Tucson)
Telescope Operator: DHuerta (LBTO)
Additional AO Support: DMiller (Tucson)
Telescope Support: JLittle (LBTO) DCox, SHooper (Tucson)
Instrument Support: JMorris (LBTO)

Summary

We completed the checkout of TCS Build 2013A spending the first 2 hours of the night exercising it with AGW1 which we didn't use last night or last week. We also exercised the build in binocular mode pinning down a few more bugs and features.

Temperature was very stable near 2-2.5 degC all night. Sky had light cirrus increasing during the night. Seeing was 0.4-0.6 arcsec FWHM.

At 5UT we switch to running AO at RFBG (w/2013A). We couldn't get any photons from CCD47 technical viewer on the DX side even though we had starlight on the pyramid (CCD39).

At 6:15 UT we switch to running AO at LFBG (w/2013A). That AO system also had problems - different ones (Issue 4484).

We finished the AGW1 realignment and checking up to the point where LUCI is required to collect transform data.

The last two hours of the night were spent taking another set of SX pointing and collimation data.

Preparation

Rotators have been repaired by installing the spare module from downtown.

Recall that this is the "new" SX pointing model from 20130124, or not.

Details

Open

Cirrus-filled sky. T=2degC.

TCS Testing and AGW startup at LFBG

01:25 Acquire Preset

01:40 IE=-83 CA=+16 EL=83

01:43 IE=-77 CA=+12 EL=77

01:45 Missed the star at EL=40 IE=-52 CA=-24 EL=40

Hold the bus! Found that PCS was using the wrong pointing model SX_Gregorian.20130110. This was because nobody restarted PCS since the new model was installed 2 days.

Load the correct ptmodel by hand = SX_Gregorian.20130124

01:51 IE=-90 CA=-4 EL=80

01:56 IE=-92 CA=0 EL=76

01:58 IE=-85 CA=-8 EL=39

Not perfect, but acceptable pointing......

02:04 Active preset -- initial collimation was good, but pinhole alignment has changed in the AGW - further complicated by this being a binary guide star.

02:05 Go to zenith and insert stowpins. Phil is checking something on LBTI. At the same time we restart PCS to load the new PCSInstrument.conf file.

Adjusting AGW1 hotspot

02:12 Active preset with Zernikes blocked. ACT0263 gs=1 EL=68
Previous Hotspot is X=225 Y=210
Cirrus in the sky helps for this. Using setHotspot left -x [X] -y [Y] to adjust the hotspot while guiding. Sending of Zernikes is disabled in GCS.

UT X Y
02:1x 230 210
02:20 225 195
02:22 225 225
02:24 225 240
02:26 225 255
02:27 240 240
02:29 255 240

Enable Zernikes for ~2 cycles

02:32 255 255
02:34 255 230

(reducing Y moves star image up on WFS)
Enable Zernikes

02:40 ACT0229 gs=1

02:45 X=255 Y=230 adopted as new hotspot and installed in LUCI_L.cfg (other places to be added later)

02:47 Offset 0 300 RADEC ABS

More offsets here

02:56 Preset with hotspot 3,7

Binocular Tests

SYNC ACQUIRE presets to ACT0229

03:10 SX Primary mirror doing some strange oscillation ?? Why?

SX IE=-96 CA=-6 DX IE=-59 CA=-131

03:23 ACTIVE preset to gs=0. Why didn't right side identify the star? Right side has lots of coma

ACTIVE ASYNC presets to gs=1

Left side looks like it is running old active optics. LUCI_IRTC_L.cfg file not updated.

03:34 offset 0 10 on SX DETXY ABS

03:35 offset 0 10 on DX DETXY ABS

03:40 SYNC offset 100 0 BC=1

03:42 SYNC offset 80 0 BC=1

03:50 SYNC preset ACTIVE to BS9126 gs=0 EL=62 (missed the right star)

03:5x ASYNC ACQUIRE on right - Left guide star got tweaked while David was adjusting IE/CA on Right. DX IE=-74 CA=-135

03:55 ASYNC ACTIVE preset on Right

03:58 ASYNC ACQUIRE on right

ptincrement 1 IE 20

ptincrement 1 IE -40 Left guide star gets dragged in the direction Right star is moving by a small fraction.

04:01 ptincrement 1 IE 40 Left guide star gets dragged in the direction Right star is moving by a small fraction.

04:04 ASYNC ACTIVE preset on Right

04:05 ASYNC ACQUIRE on Left IE=-96 CA=-6 not centered

04:07 ptincrement 0 IE 40 -- tweaks Right guide star in same direction about 2 arcsec (Issue 4514)

04:08 ptincrement 0 IE -40

04:09 ptincrement 0 CA -40

04:10 ptincrement 0 CA -40

04:12 newpoint BS9126 TelM=ACTIVE gs=11 BC=1

04.3 Star guiding on SX systematically +x from hotspot as seen on guideplot WHY?

04:28 newpoint BS9126 TelM=ACTIVE gs=16 BC=1

04:34 newpoint ACT0229 TelM=ACQUIRE gs=0 BC=1 and recenter

SX IE=-100 CA=13 DX IE=-54 CA=-132

04:49 Doug M takes the helm, but not the hot seat.

Adaptive Optics Tests as IRTC @ RFBG

Seeing (w/o adaptive optics) is 0.42 arcsec FWHM on guider. DIMM says 0.6.

05:12 ADAPTIVE_ACE preset - No star found in TV frame.

05:46 Preset to Pollux -- We have starlight on the DX pyramid, but not on the technical viewer.

05:55 Shut down DX AO and restart.

06:00 ADAPTIVE_ACE preset - No joy on CCD47.

06:13 Authorize LUCI1 @ LFBG

Adaptive Optics Tests as XXXX @ LFBG

06:30 ADAPTIVE_ACE preset "PresetAO[136] failed - RETRY: No settings for LUCIFER - ACE-A"

06:35 Switch to IRTC @ LFBG

06:40 ADAPTIVE_ACE preset - no star - IRTCF collimation model hadn't been updated.

06:45 cp -p SXPMLUCIFERCollimation.dat SXPMIRTCFCollimation.dat
06:45 cp -p SXSMLUCIFERCollimation.dat SXSMIRTCFCollimation.dat

06:49 ADAPTIVE_ACE_ACTIVE preset - camera lens centering did not converge? or was CCD39 hung?

06:59 ADAPTIVE_ACE_ACTIVE preset w/ different position angle - same error

07:07 ADAPTIVE_ACE_ACTIVE preset - same error

AGW1 Alignment

07:14 Authorize LUCI1@LFBG

Measure and Adjust Rotator Center on AGW1 Guider

Data in /home/aoeng/data/20130202/

07:32:50 image 0002

[telescope@agw1 ~]$ getxy -u 1
Guide probe position
	x: -0.015000
	y: 612.500000

[telescope@agw1 ~]$ getxy -u 1 -o 16
Guide probe position
	x: -1.000000
	y: 1012152.000000

Want to match new hotspot of 255,230 which is 127.5,155 in binned coordinates.

07:35 guideimage 0003 measured X=120.5 Y=117 r=47.13

07:50 subtracted 68 counts from ROT homeoffset for AGW1. 35578 --> 35510 updated x_0, y_0, x_guider, y_guider

07:51 stop and restart GCSs, AGW1, oacserver

08:01 guideimage 0004 hung. Had to kill and restart GCS

08:06 guideimage 0006 X=137.5 Y=116 - moved to far

08:15 added 38 counts from ROT homeoffset for AGW1. 35510 --> 35548, updated lim_f, lim_n, took out some spaces.

08:17 stop and restart GCSs, AGW1, oacserver

08:22 ACTIVE preset

08:23 ACQUIRE preset

08:25 ACTIVE preset

[telescope@agw1 etc]$ getxy -u 1
Guide probe position
	x: 0.015000
	y: 612.499000

[telescope@agw1 etc]$ getxy -u 1 -o 16
Guide probe position
	x: 1.000000
	y: 1012150.000000

08:27 guideimage 0008 X=128.25 Y=115.5

Done with this adjustment.

Checking Pupil Wobble on AGW1

Preset ACTIVE on-axis BS9126

08:38 stop WFSing to make it a guide preset

08:40 Run pupil_collect to check pupil wobble and M3 alignment

Guiding delay is a bit too short for 60 degree offsets.

Angle Xpup Ypup
0 387.6 206.9
60 388.4 208.2
120 387.3 209.3
180 385.1 208.8
240 385.8 207.2
300 385.1 206.6

08:50 Done adopt pupil center 386.0, 207.5 Put in LUCI_L.cfg, and comment out WFS_hotspot_sext_x and y

08:57 ACTIVE preset on-axis

Checking off-axis pupil position

09:04 Preset off-axis

09:13 uncomment hotspot_sext_x and y in LUCI_L.cfg - those parameters still do something.

Studying the bad column of spots at the right edge of AGw1 wfs.

09:23 Preset

Collimation Zeropoints

09:30 In SXSMLUCIFERCollimation.dat changed from:
 -6.175530e+00 -2.807068e+00 +6.745767e-01 +5.289454e+02 +2.749935e+02 -1.742340e-01  Offset collimation coefficient 
To:
 -5.405530e+00 -2.537068e+00 +6.745767e-01 +5.009454e+02 +3.549935e+02 -1.742340e-01  Offset collimation coefficient
These were the active optics offsets from tonight, and these corrections were 80% of what DLM measured from the previous pointing/collimation data.

Note: Later in the night, it looked like we should have used 60% of Doug's offsets rather than 80%.

09:40 WFS_SubDir changed from 20120816 to 20121213 in LUCI_L.cfg (masking bad column on right for LUCI).

Pointing Zeropoints

09:50 Active preset BS9139

Adopt IE=-83 CA=-7 for SX edit SX_Gregorian.20130124.ptmod

09:56 Restart PCS

Field Data at LFBG at BS9155

09:59 Seeing on guider at EL=58 is down to 0.35 arcsec. These conditions are too good to spend on pointing data, so we'll take a set of field data at LFBG. Thicker clouds are approaching from the west.

10:02 Active preset to BS9155 at EL=35

Data in /home/aoeng/data/20130202/field_01/

10:11 field_collect with BS9155_field.list and limit=200

10:35 0.5 arcsec on guider at EL=40 -- Sweet. From the numbers scrolling by in IDL, this is good data, and there is no significant binodal astigmatism.

11:13 field data finished without any interruption or restarting

GStar wfsc #
0 322-323
13 324-326
0 327-329
15 330-332
0 333-335
9 336-338
0 339-341
11 342-344
0 345-346
12 347-349
0 350-352
5 353-355
0 363-358
6 364-366
0 367
4 368-370
0 374-373
7 375-377
0 378-379
8 380-382
0 383-384

SX Pointing Data @ LFBG

11:13 Start pointing log: /home/telescope/PointingLogs/20130202_SX.log

Data lives in /home/aoeng/data/20130202/pointing/

11:15 ptmodel_collect (default LIMIT=500)

Note: This pointing data is very slow because we are collimating out the astigmatism at each change in elevation. That is taking 3-5 iterations per star. For the first 12 stars, I assume the time is useful as a check of the collimation model.

11:30 missed the star on ACT0350 at EL=61 AZ=-55 adjust IE=-80 CA=-7

11:32 ptmodel_collect, LIMIT=1000, START=4

11:37 WT10_288 is a binary star. Perhaps not good coordinates.

11:41 In LUCI_L.cfg reduced WFS_exposure from 30000 to 10000 to save collimation time. It took effect on the fly.

Other things that could be wrong with pointing:
  • Refraction: LUCI is set for lambda_target=1.65 (should match the guider)
  • We are using off-axis guide stars, and the transformation might not be up to date - AGW1 has been dismounted.

11:55 ACT0249 was binary - bad coordinates.

WT10_266 could have picked the wrong star

12:04 missed WT10_311 at EL=82 (number 16)

12:06 ptmodel_collect, LIMIT=1000, START=16

Sat Feb 2 12:37:34.277 2013 Preset alert for left LUCIFER mode ACTIVE: sx.adsc: [moveAbs.sendString] Command not allowed unless buffer open

12:39 ptmodel_collect, LIMIT=1000, START=32

12:44 Star not found due to clouds.......skipped grid 34

12:45 ptmodel_collect, LIMIT=1000, START=35

WT10_190 didn't have guide stars available.

13:20 Star not found due to clouds, skipped grid 54.

13:30 Star not found due to clouds and twilight.

Star Name Target Star GStar WFSC Image
1 WT10_365 2 2 left_wfscimage000392.fits
2 ACT0411 1 1 left_wfscimage000397.fits
3 WT10_404 4 4 left_wfscimage000400.fits
4 ACT0350 2 2 left_wfscimage000403.fits
5 WT10_288 7 7 left_wfscimage000407.fits
6 ACT0203 1 1 left_wfscimage000408.fits
7 WT10_241 1 1 left_wfscimage000408.fits
8 WT10_135 1 1 left_wfscimage000409.fits
9 WT10_192 1 1 left_wfscimage000411.fits
10 ACT0279 3 3 left_wfscimage000414.fits
11 ACT0249 13 13 left_wfscimage000419.fits
12 ACT0330 10 10 left_wfscimage000422.fits
13 WT10_343 1 1 left_wfscimage000424.fits
14 ACT0283 2 2 left_wfscimage000425.fits
15 WT10_266 2 2 left_wfscimage000427.fits
16 WT10_311 7 7 left_wfscimage000431.fits
17 ACT0209 1 1 left_wfscimage000434.fits
18 WT10_122 15 15 left_wfscimage000438.fits
20 WT10_294 1 1 left_wfscimage000439.fits
21 ACT0258 2 2 left_wfscimage000441.fits
22 WT10_220 3 3 left_wfscimage000444.fits
23 WT10_339 2 2 left_wfscimage000449.fits
24 WT10_407 3 3 left_wfscimage000454.fits
25 WT10_406 3 3 left_wfscimage000457.fits
26 WT10_378 1 1 left_wfscimage000459.fits
27 WT10_389 11 11 left_wfscimage000460.fits
28 ACT0396 6 6 left_wfscimage000461.fits
32 ACT0409 2 2 left_wfscimage000463.fits
33 WT10_365 2 2 left_wfscimage000466.fits
35 WT10_350 1 1 left_wfscimage000468.fits
36 ACT0363 2 2 left_wfscimage000469.fits
37 ACT0324 1 1 left_wfscimage000471.fits
38 WT10_309 1 1 left_wfscimage000472.fits
39 WT10_258 3 3 left_wfscimage000475.fits
40 WT10_186 2 2 left_wfscimage000476.fits
41 ACT0188 2 2 left_wfscimage000476.fits
42 WT10_262 1 1 left_wfscimage000480.fits
43 ACT0209 1 1 left_wfscimage000482.fits
44 WT10_175 2 2 left_wfscimage000484.fits
45 ACT0141 2 2 left_wfscimage000487.fits
46 ACT0177 1 1 left_wfscimage000489.fits
47 WT10_266 2 2 left_wfscimage000492.fits
48 ACT0298 1 1 left_wfscimage000494.fits
49 ACT0315 1 1 left_wfscimage000495.fits
50 WT10_302 1 1 left_wfscimage000499.fits
51 ACT0358 5 5 left_wfscimage000500.fits
52 WT10_347 2 2 left_wfscimage000501.fits
55 ACT0343 4 4 left_wfscimage000504.fits
56 ACT0328 1 1 left_wfscimage000506.fits
57 WT10_315 2 2 left_wfscimage000509.fits
58 ACT0283 2 2 left_wfscimage000510.fits
59 ACT0232 1 1 left_wfscimage000511.fits
60 WT10_196 2 2 left_wfscimage000513.fits

Close

13:32 Close for impending dawn and clouds.

-- JohnHill - 02 Feb 2013
Topic attachments
I Attachment Action Size Date Who Comment
daq.20130202telescope.jpgjpg daq.20130202telescope.jpg manage 587 K 05 Feb 2013 - 21:38 JohnHill Plot of Telescope Temperatures
Topic revision: r7 - 19 Feb 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback